OpenStudy (shadow13):

i need help please If a and b are odd integers, then a • b is an odd integer. Which is the inverse of the statement? If a and b are not odd integers, then a • b is not an odd integer. If a • b is an odd integer, then a and b are odd integers. a and b are odd integers a • b is an odd integer

OpenStudy (michele_laino):

hint: if a and b are odd integers, then exists two integers, say m, and n, such that: \[\begin{gathered} a = 2n + 1 \hfill \\ b = 2m + 1 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so we can write: \[a \cdot b = 4nm + 2n + 2m + 1\]